OpenStudy (anonymous):

|5x-2| greater than or equal to 8.....I got x greater than or equal to 1.6 or x less than or equal to 2. How do I graph this?

OpenStudy (anonymous):

\[|5x-2|\geq8\]

OpenStudy (anonymous):

\[5x-2\geq8\] or \[-(5x-2)\geq8\]

OpenStudy (anonymous):

\[-(5x-2)\geq8\] divide by -1 \[(5x-2)\leq-8\]

OpenStudy (anonymous):

First of all draw two lines y=(5x-2) and y=-(5x-2)

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

1) Draw two lines y=(5x-2) and y=-(5x-2) 2) Sub in some random number (x) and shade in

OpenStudy (anonymous):

i need to graph the solutions x is greater than or equal to 2 or x is less than or equal to -1.2....please help